### v1

@@ Abstract
After pre-training on extensive image-text pairs, Contrastive Language-Image Pre-training (CLIP) demonstrates promising performance on a wide variety of benchmarks. However, a substantial volume of non-paired data, such as multimodal interleaved documents, remains underutilized for vision-language representation learning. To fully leverage these unpaired documents, we initially establish a Real-World Data Extraction pipeline to extract high-quality images and texts. Then we design a hierarchical retrieval method to efficiently associate each image with multiple semantically relevant realistic texts. To further enhance fine-grained visual information, we propose an image semantic augmented generation module for synthetic text production. Furthermore, we employ a semantic balance sampling strategy to improve dataset diversity, enabling better learning of long-tail concepts. Based on these innovations, we construct RealSyn, a dataset combining realistic and synthetic texts, available in three scales: 15M, 30M, and 100M. Extensive experiments demonstrate that RealSyn effectively advances vision-language representation learning and exhibits strong scalability. Models pre-trained on RealSyn achieve state-of-the-art performance on multiple downstream tasks. To facilitate future research, the RealSyn dataset and pre-trained model weights are released at https://github.com/deepglint/RealSyn. footnotetext: * Equal contribution. footnotetext: $\dagger$ Corresponding author.
RealSyn: An Effective and Scalable Multimodal Interleaved Document Transformation Paradigm
[…]
1 Introduction
The rapid proliferation of mobile networks and social platforms has led to exponential growth in large-scale data, offering a robust foundation for vision-language representation learning Guo et al. (2019); Baltrušaitis et al. (2018); Guo et al. (2024). By predicting the correspondence between images and description texts, Contrastive Language-Image Pre-training (CLIP) Radford et al. (2021) pretrains separate uni-modal encoders and achieves excellent performance across a range of downstream tasks Tian et al. (2020); He et al. (2020); Chen et al. (2020); Chopra et al. (2005); Shao et al. (2024); Sun et al. (2024); Luo et al. (2024). Notably, the representational capacity of such models improves significantly with the scale of the training dataset Li et al. (2024b). However, the availability of paired data in the real world is limited.
Figure 1: Multimodal interleaved documents are unsuitable for vision-language representation learning. We construct distinct image-text pairs from such documents via retrieval and generation.
In recent years, the release of large-scale image-text pair datasets Schuhmann et al. (2021, 2022); Thomee et al. (2016); Byeon et al. (2022); Gadre et al. (2024) has significantly advanced vision-language representation learning. Due to download failures and non-English captions in the YFCC Thomee et al. (2016) dataset, DeCLIP Li et al. (2022) reprocesses a new version of the YFCC15M dataset. LAION400M Schuhmann et al. (2021) provides 400 million image-text pairs collected from the web and filtered using CLIP similarity. COYO700M Byeon et al. (2022) offers 747 million high-quality pairs curated through advanced filtering strategies, including CLIP similarity, watermark detection, and aesthetic scoring. Despite these advancements, a substantial amount of non-paired data, particularly interleaved image-text documents Zhu et al. (2024); Laurençon et al. (2024); Li et al. (2025) containing multiple images and text paragraphs without explicit correspondence, remains incompatible with conventional vision-language representation learning methods.
Recent studies Fan et al. (2024); Yu et al. (2024); Zheng et al. (2024a); Lai et al. (2024) aim to enhance the quality of open-source image-text datasets using existing models. For instance, LaCLIP Fan et al. (2024) leverages large language models (LLMs) to rewrite raw captions for better alignment with images. CapsFusion Yu et al. (2024) fine-tunes an LLM using a ChatGPT-generated instruction dataset to mitigate hallucination issues. However, the diversity and distribution of synthetic data generated by these methods are inherently constrained by the limitations of the underlying generative models.
Figure 2: The Real-World Data Extraction pipeline to extract high-quality images and texts from interleaved image-text documents.
In this paper, we explore two fundamental questions: 1) How to utilize multimodal interleaved documents for vision-language representation learning. 2) How to effectively leverage both realistic and synthetic texts to enhance representation performance. To this end, as shown in Fig. 1, we first establish a Real-World Data Extraction pipeline to extract high-quality images and texts. Then we design a hierarchical retrieval method to efficiently associate each image with multiple semantically relevant texts. To enhance fine-grained image understanding, we propose a visual semantic augmented generation module for synthetic text production. Furthermore, we employ a semantic balance sampling strategy to improve dataset diversity, enabling better learning of long-tail concepts. Based on these innovations, we construct the RealSyn dataset, which integrates both realistic and synthetic texts with three sizes: 15M, 30M, and 100M. Comprehensive experimental results show that RealSyn is effective for vision-language representation learning and exhibits excellent scalability. The main contributions of this paper are summarized as follows:
@@ 1 Introduction
  • We propose an effective and scalable multimodal interleaved document transformation paradigm for vision-language representation learning.
@@
  • We conduct extensive experiments and demonstrate the effectiveness and scalability of our proposed RealSyn dataset.
2 Related Work
@@
Large-Scale Pre-training Dataset. In recent years, several large-scale image-text datasets Chen et al. (2015); Clark and Gardner (2017); Goyal et al. (2017); Byeon et al. (2022); Li et al. (2024a) collected from the Internet have been released. The YFCC100M Thomee et al. (2016) dataset provides a comprehensive overview of the evolution of photo and video documentation and sharing from the inception of Flickr in 2004 until early 2014. Due to download failures and non-English captions, DeCLIP Li et al. (2022) reprocesses a new version of the YFCC15M dataset. Additionally, the LAION400M Schuhmann et al. (2021) dataset contains 400 million image-text pairs collected from Common Crawl and widely used in vision-language pre-training. Recent advancements have also introduced several large-scale interleaved image-text document datasets Li et al. (2025); Zhu et al. (2024); Laurençon et al. (2024). The OBELICS Laurençon et al. (2024) dataset uses a comprehensive filtering strategy and includes 141 million web pages, 353 million associated images, and 115 billion text tokens extracted from Common Crawl. However, due to data format constraints and training inefficiencies, interleaved image-text documents are currently unsuitable for vision-language representation learning.
@@
Vision Language Pre-training. As a pioneering work in visual language pre-training, CLIP has attracted extensive attention due to its powerful zero-shot recognition and exceptional transfer learning performance Wang and Kang (2024); Tang et al. (2024); Shao et al. (2024); Martin et al. (2024). Inspired by CLIP, numerous visual-language pre-training works have been published in recent years Mu et al. (2022); Li et al. (2022); Yang et al. (2023). SLIP Mu et al. (2022) enhances performance by combining self-supervised learning with CLIP pre-training. DeCLIP Li et al. (2022) improves pre-training efficiency by integrating multi-view supervision across modalities and nearest-neighbor supervision from similar pairs. To mitigate the influence of noisy data, ALIP Yang et al. (2023) introduces a gating mechanism to dynamically allocate weights to samples. However, all these methods rely on large-scale image-text pairs crawled from the Internet.
@@
Synthetic Captions. Recent works Yang et al. (2023); Yu et al. (2024); Chen et al. (2024) indicate that image-text pairs obtained from websites contain intrinsic noise, which directly impacts the effectiveness of vision-language pre-training. To enhance data diversity, LaCLIP Fan et al. (2024) uses the in-context learning capability of large language models to rewrite text descriptions associated with each image. CapsFusion Yu et al. (2024) employs large language models to refine information from web-based image-text pairs and synthetic captions, improving the quality of multimodal pre-training data. However, the quality and diversity of synthetic data are inherently limited by the capabilities of the generative model.
3 RealSyn Dataset
3.1 Real-World Data Extraction
@@
To transform interleaved image-text documents for vision-language representation learning, we establish a Real-World Data Extraction pipeline (Fig. 2) to extract high-quality images and texts. This pipeline consists of three steps: Data Extraction, Image Filtration, and Sentence Filtration.
@@
Data Extraction. We employ 118M interleaved image-text documents from the OBELICS Laurençon et al. (2024) as the primary data source. All images are extracted and stored in a dedicated image database, while sentences are segmented using the Natural Language Toolkit (NLTK) Bird and Loper (2004) and stored in a separate sentence database. This process yields 336M images and 2.13B sentences from the interleaved documents.
@@
Image Filtration. After extracting 336M images, we apply a two-stage filtering process to ensure data quality and reduce redundancy. First, we discard images that meet any of the following criteria: 1) the shorter dimension is fewer than 100 pixels, or 2) the aspect ratio exceeds 3 or is below 1/3. This step removes 51M low-quality images. Next, following CLIP-CID Yang et al. (2025), we use the EVA02-CLIP E/14-plus model Sun et al. (2023) to extract image embeddings and apply the Union-Find algorithm Tarjan (1975) to eliminate perceptually and semantically redundant images. This step removes an additional 87M images, resulting in a refined dataset of 198M high-quality images.
@@
Sentence Filtration. After extracting 2.13B sentences from interleaved image-text documents, we conduct rigorous filtering based on quality, semantics, and redundancy. Initially, we eliminate sentences based on the following criteria: 1) presence of emojis or URLs; 2) sentences containing fewer than 3 or more than 81 words; and 3) following CAT Radenovic et al. (2023), we retain samples with at least C1 caption complexity and incorporating an action. This phase reduces the corpus size from 2.13B to 1.82B sentences. Then we apply semantic filtering to the remaining sentences, eliminating those with minimal information assessed through information entropy:
$\theta(x)=-\sum_{i=1}^{n}p(x_{i})\log p(x_{i}),$ (1)
@@
where $n$ denotes the number of words in a sentence, $x_{i}$ represents the $i$-th words in the sentences $x$, and $p(x_{i})$ is the probability of the word $x_{i}$ in the entire corpus. Based on human cognition principles and empirical experience, we filter out sentences with a score below 0.3. To further refine the corpus by removing difficult or ambiguous sentences, we use GTP2-large Radford et al. (2019) to calculate the perplexity score $\mathcal{PPL}$ for each sentence:
$\mathcal{PPL}(x)=\exp\left\{-\frac{1}{t}\sum_{i=1}^{t}\log p_{\theta}(x_{i}\mid x_{<i})\right\},$ (2)
@@
where $t$ represents the token number of the sentence, and $p_{\theta}(x_{i}\mid x_{<i})$ is the likelihood of the $i$-th token given the previous tokens. We retain sentences with perplexity scores between 30 and 200. The overall semantics filtering reduces the corpus to 1.16B sentences. In the final stage, similar to redundancy image filtering, we perform both perceptual and semantic deduplication of sentences. This process results in a refined corpus of 0.84B sentences that include extensive real-world knowledge.
@@
Figure 3: The architecture of our proposed framework, which constructs distinct image-text pairs from real-world data extracted from interleaved documents via retrieval and generation.
3.2 Retrieval and Generation Framework
After extracting high-quality images and sentences from documents, we propose an efficient and scalable framework to retrieve multiple semantically relevant texts for each image and leverage large language models to integrate retrieved realistic text with fine-grained visual information and generate synthetic text. As shown in Fig. 3, the architecture of our framework primarily consists of three components: Text Semantic Clustering, Hierarchical Retrieval, and Image Semantic Augmented Generation.
@@ 3.2 Retrieval and Generation Framework
Table 1: Linear probe on 20 downstream datasets. Pre-training ViT-B/32 on RealSyn achieves 1.3%-6.9% average performance improvement.
@@
Table 2: Zero-shot transfer on 20 downstream datasets. Pre-training ViT-B/32 on RealSyn achieves 2.3%-14.3% average performance improvement.

### v3

@@ Abstract.
After pre-training on extensive image-text pairs, Contrastive Language-Image Pre-training (CLIP) demonstrates promising performance on a wide variety of benchmarks. However, a substantial volume of multimodal interleaved documents remains underutilized for contrastive vision-language representation learning. To fully leverage these unpaired documents, we initially establish a Real-World Data Extraction pipeline to extract high-quality images and texts. Then we design a hierarchical retrieval method to efficiently associate each image with multiple semantically relevant realistic texts. To further enhance fine-grained visual information, we propose an image semantic augmented generation module for synthetic text production. Furthermore, we employ a semantic balance sampling strategy to improve dataset diversity, enabling better learning of long-tail concepts. Based on these innovations, we construct RealSyn, a dataset combining realistic and synthetic texts, available in three scales: 15M, 30M, and 100M. We compare our dataset with other widely used datasets of equivalent scale for CLIP training. Models pre-trained on RealSyn consistently achieve state-of-the-art performance across various downstream tasks, including linear probe, zero-shot transfer, zero-shot robustness, and zero-shot retrieval. Furthermore, extensive experiments confirm that RealSyn significantly enhances contrastive vision-language representation learning and demonstrates robust scalability.
@@
large-scale image-text dataset, vision-language representation learning, multimodal interleaved document
* Equal Contribution
$\dagger$ Corresponding Author
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755098isbn: 979-8-4007-2035-2/2025/10
1. Introduction
The rapid proliferation of mobile networks and social platforms has led to exponential growth in large-scale data, offering a robust foundation for contrastive vision-language representation learning (Guo et al., 2019; Baltrušaitis et al., 2018; Guo et al., 2024a; Zhu et al., 2024a; Yao et al., 2024; Zheng et al., 2024a; Guo et al., 2024b). By predicting the correspondence between images and description texts, Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) pre-trains separate uni-modal encoders and achieves excellent performance across a range of downstream tasks, including image captioning (Li et al., 2023; Mokady et al., 2021), object detection (Wu et al., 2023; Lin and Gong, 2023), and semantic segmentation (He et al., 2023; Zhou et al., 2023; Lin et al., 2023). Notably, the representational capacity of such models improves significantly with the scale of the training dataset (Li et al., 2024a). Designing a new data paradigm to expand the scale and diversity of image-text datasets, thereby continuously improving the capabilities of pre-trained vision-language representations, represents a significant and challenging issue.
Figure 1. Multimodal interleaved documents are unsuitable for CLIP training. We construct distinct image-text pairs from such documents via retrieval and generation.
Figure 2. The Real-World Data Extraction pipeline to extract high-quality images and texts from interleaved image-text documents.
In recent years, the introduction of large-scale image-text pair datasets (Schuhmann et al., 2021, 2022; Thomee et al., 2016; Byeon et al., 2022; Gadre et al., 2024) has substantially advanced vision-language representation learning. Due to download failures and non-English captions in the YFCC (Thomee et al., 2016) dataset, DeCLIP (Li et al., 2022) reprocesses a new version of the YFCC15M dataset. LAION400M (Schuhmann et al., 2021) provides 400 million image-text pairs collected from the web and filtered using CLIP similarity. COYO700M (Byeon et al., 2022) offers 747 million high-quality pairs curated through advanced filtering strategies, including CLIP similarity, watermark detection, and aesthetic scoring. Despite these advancements, a substantial amount of non-paired data, particularly interleaved image-text documents (Zhu et al., 2024b; Laurençon et al., 2024; Li et al., 2025) containing multiple images and text paragraphs without explicit correspondence, remains incompatible with conventional vision-language representation learning methods.
Recent studies (Fan et al., 2024; Yu et al., 2024; Zheng et al., 2024b; Lai et al., 2024) aim to enhance the quality of open-source image-text datasets using existing models. For instance, LaCLIP (Fan et al., 2024) leverages large language models (LLMs) to rewrite raw captions for better alignment with images. CapsFusion (Yu et al., 2024) fine-tunes an LLM using a ChatGPT-generated instruction dataset to mitigate hallucination issues. DreamLIP (Zheng et al., 2024b) re-captions 30 million images with detailed descriptions using a pre-trained large multimodal model. However, these methods primarily enhance data quality by focusing on synthetic data, thereby overlooking the significance of real-world data. Moreover, the diversity and distribution of synthetic captions produced by these approaches are inherently restricted by the limitations of the underlying generative models.
In this paper, we explore two fundamental questions: 1) How to utilize multimodal interleaved documents for CLIP training. 2) How to effectively leverage both realistic and synthetic texts to enhance CLIP performance. As depicted in Figure 1, we initially develop a Real-World Data Extraction pipeline to extract high-quality images and texts from real-world multimodal interleaved documents. Subsequently, we design a hierarchical retrieval method to efficiently associate each image with multiple semantically relevant texts. To improve fine-grained image understanding, we introduce a visual semantic augmented generation module for synthetic text production. Additionally, we implement a semantic balance sampling strategy to enhance dataset diversity and facilitate the learning of long-tail concepts. Leveraging these innovations, we construct the RealSyn dataset, which incorporates both realistic and synthetic texts in three sizes: 15M, 30M, and 100M. We evaluate the RealSyn dataset against other widely used datasets of similar scale for CLIP training. Models pre-trained on RealSyn consistently demonstrate state-of-the-art performance on various downstream tasks, including linear probe, zero-shot transfer, zero-shot robustness, and zero-shot retrieval. Extensive experiments validate that RealSyn is effective for contrastive vision-language representation learning and shows excellent scalability. The main contributions of this paper are summarized as follows:
  • We propose an effective and scalable multimodal interleaved document transformation paradigm for contrastive vision-language representation learning.
@@ 1. Introduction
  • We conduct extensive experiments and demonstrate the effectiveness and scalability of our proposed RealSyn dataset for CLIP training.
2. Related Work
@@
Large-Scale Pre-training Dataset. In recent years, several large-scale image-text datasets (Chen et al., 2015; Clark and Gardner, 2017; Goyal et al., 2017; Byeon et al., 2022; Li et al., 2024b) collected from the Internet have been released. The YFCC100M (Thomee et al., 2016) dataset provides a comprehensive overview of the evolution of photo and video documentation and sharing from the inception of Flickr in 2004 until early 2014. Due to download failures and non-English captions, DeCLIP (Li et al., 2022) reprocesses a new version of the YFCC15M dataset. Additionally, the LAION400M (Schuhmann et al., 2021) dataset contains 400 million image-text pairs collected from Common Crawl and widely used in vision-language pre-training. Recent advancements have also introduced several large-scale interleaved image-text document datasets (Li et al., 2025; Zhu et al., 2024b; Laurençon et al., 2024). The OBELICS (Laurençon et al., 2024) dataset uses a comprehensive filtering strategy and includes 141 million web pages, 353 million associated images, and 115 billion text tokens extracted from Common Crawl. However, due to data format constraints and training inefficiencies, interleaved image-text documents are currently unsuitable for CLIP training.
@@
Vision Language Pre-training. As a pioneering work in visual language pre-training, CLIP has attracted extensive attention due to its powerful zero-shot recognition and exceptional transfer learning performance (Wang and Kang, 2024; Tang et al., 2024; Shao et al., 2024; Martin et al., 2024; Hu et al., 2025; Gu et al., 2024). Inspired by CLIP, numerous visual-language pre-training works have been published in recent years (Mu et al., 2022; Li et al., 2022; Yang et al., 2023). SLIP (Mu et al., 2022) enhances performance by combining self-supervised learning with CLIP pre-training. DeCLIP (Li et al., 2022) increases pre-training efficiency by integrating multi-view supervision across modalities and nearest-neighbor supervision from similar pairs. To mitigate the influence of noisy data, ALIP (Yang et al., 2023) introduces a gating mechanism that dynamically allocates weights to samples. Despite their advancements, these methods primarily depend on large-scale image-text pairs derived from the Internet. Recent studies (Li et al., 2024a; Wang et al., 2025) demonstrate that the capabilities of CLIP enhance with the expansion of high-quality image-text datasets. Given the extensive use of internet-derived image-text data in existing datasets, there is a pressing need to develop a new data construction paradigm to further expand the scale of high-quality image-text data.
@@
Synthetic Captions. Recent works (Yang et al., 2023; Yu et al., 2024; Chen et al., 2024) indicate that image-text pairs obtained from websites contain intrinsic noise, which directly impacts the effectiveness of vision-language pre-training. To enhance the quality of existing datasets, LaCLIP (Fan et al., 2024) uses the in-context learning capability of large language models to rewrite text descriptions associated with each image. CapsFusion (Yu et al., 2024) employs large language models to refine information from web-based image-text pairs and synthetic captions, improving the quality of multimodal pre-training data. Similarly, DreamLIP (Zheng et al., 2024b) generates detailed descriptions for 30 million images using a pretrained large multimodal model. Nevertheless, these methods predominantly focus on synthetic data enhancement, neglecting the importance of real-world data. Furthermore, the diversity and distribution of synthetic captions generated by these methods are intrinsically constrained by the capabilities of the generative models employed.
3. RealSyn Dataset
3.1. Real-World Data Extraction
@@
To transform interleaved image-text documents for vision-language representation learning, we establish a Real-World Data Extraction pipeline (Figure 2) to extract high-quality images and texts. This pipeline consists of three steps: Data Extraction, Image Filtration, and Sentence Filtration.
@@
Data Extraction. We employ 118M interleaved image-text documents from the OBELICS (Laurençon et al., 2024) as the primary data source. All images are extracted and stored in a dedicated image database, while sentences are segmented using the Natural Language Toolkit (NLTK) (Bird and Loper, 2004) and stored in a separate sentence database. This process yields 336M images and 2.13B sentences from the interleaved documents.
@@
Image Filtration. After extracting 336M images, we apply a two-stage filtering process to ensure data quality and reduce redundancy. First, we discard images that meet any of the following criteria: 1) the shorter dimension is fewer than 100 pixels, or 2) the aspect ratio exceeds 3 or is below 1/3. This step removes 51M low-quality images. Next, following CLIP-CID (Yang et al., 2025), we use the EVA02-CLIP E/14-plus model (Sun et al., 2023) to extract image embeddings and apply the Union-Find algorithm (Tarjan, 1975) to eliminate perceptually and semantically redundant images. This step removes an additional 87M images, resulting in a refined dataset of 198M high-quality images.
@@
Sentence Filtration. After extracting 2.13B sentences from interleaved image-text documents, we conduct rigorous filtering based on quality, semantics, and redundancy. Initially, we eliminate sentences based on the following criteria: 1) presence of emojis or URLs; 2) sentences containing fewer than 3 or more than 81 words; and 3) following CAT (Radenovic et al., 2023), we retain samples with at least C1 caption complexity and incorporating an action. This phase reduces the corpus size from 2.13B to 1.82B sentences. Then we apply semantic filtering to the remaining sentences, eliminating those with minimal information assessed through information entropy:
(1) $\theta(x)=-\sum_{i=1}^{n}p(x_{i})\log p(x_{i}),$
@@
where $n$ denotes the number of words in a sentence, $x_{i}$ represents the $i$-th words in the sentences $x$, and $p(x_{i})$ is the probability of the word $x_{i}$ in the entire corpus. Based on human cognition principles and empirical experience, we filter out sentences with a score below 0.3. To further refine the corpus by removing difficult or ambiguous sentences, we use GTP2-large (Radford et al., 2019) to calculate the perplexity score $\mathcal{PPL}$ for each sentence:
(2) $\mathcal{PPL}(x)=\exp\left\{-\frac{1}{t}\sum_{i=1}^{t}\log p_{\theta}(x_{i}\mid x_{<i})\right\},$
@@
where $t$ represents the token number of the sentence, and $p_{\theta}(x_{i}\mid x_{<i})$ is the likelihood of the $i$-th token given the previous tokens. We engage three human experts to determine the minimum and maximum values of the perplexity interval for high-quality sentences by comparing sentences of varying perplexity levels. Sentences within the average minimum (30) and maximum perplexity (200) intervals are retained. The overall semantics filtering reduces the corpus to 1.16B sentences. In the final stage, similar to redundancy image filtering, we perform both perceptual and semantic deduplication of sentences. This process results in a refined corpus of 0.84B sentences that include extensive real-world knowledge.
@@
Figure 3. The architecture of our proposed framework, which constructs distinct image-text pairs from real-world data extracted from interleaved documents via retrieval and generation.
Table 1. Linear probe on 20 downstream datasets. Pre-training ViT-B/32 on RealSyn achieves 1.3%-6.9% average performance improvement.
@@
Table 2. Zero-shot transfer on 20 downstream datasets. Pre-training ViT-B/32 on RealSyn achieves 2.3%-14.3% average performance improvement.
@@
3.2. Retrieval and Generation Framework
After extracting high-quality images and sentences from documents, we propose an efficient and scalable framework to retrieve multiple semantically relevant texts for each image and leverage large language models to integrate retrieved realistic text with fine-grained visual information and generate synthetic text. As shown in Figure 3, the architecture of our framework primarily consists of three components: Text Semantic Clustering, Hierarchical Retrieval, and Image Semantic Augmented Generation.